### v4

# King’s College London, U.K

Abstract
We demonstrate that Assembly Theory (AT) is reliant on a method based upon statistical compression that belongs to the LZ family of compression algorithms, family which comprises popular algorithms used in ZIP, GZIP, or JPEG. Such popular algorithms have been shown to empirically reproduce the results of AT that were reported before in successful applications to separating organic from non-organic molecules and in the context of the study of selection and evolution. In particular, assembly index is a LZ compression scheme whose statistical compressibility is bounded by entropy or other traditional LZ compression schemes, such as LZ77, LZ78, or LZW. In addition, we demonstrate that AT, its pathway complexity, the assembly index, and the assembly number define compression schemes that are subsumed into the theory of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity. Due to AT’s current lack of logical consistency in defining causality for non-stochastic processes and lack of the empirical evidence that it outperforms other complexity measures found in the literature, we conclude that the assembly index and the assembly number do not lead to an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by (abiotic or Darwinian) selection and evolution.
keywords:
assembly theory, assembly index, pathway complexity, assembly number, algorithmic complexity, algorithmic probability, CTM, BDM, lossless compression, LZ compression, LZ77, LZ78, LZW, Lempel–Ziv–Welch, logical depth, selection, evolution.
1 Introduction
Assembly Theory (AT) has recently garnered significant attention in some academic circles and in the scientific media. Responding to how open-ended forms can emerge from matter without a blueprint, AT purports to explain and quantify selection and evolution. The central claim of Assembly Theory advanced in [1, 2] is that objects with a high assembly index “are very unlikely to form abiotically”. While this has been contested in [3], we will demonstrate the connections of Assembly Theory and its assembly index to well-established theories that the authors claim are different and unrelated to Assembly Theory [4]. We will show that the claims advanced by its authors that Assembly Theory unifies life and biology with physics [2, 5] rely on a circular argument and the use of a popular compression algorithm, with no empirical or logical support to establish deeper connections to selection and evolution not already made before [6, 7]. We demonstrate how Assembly Theory is an approximation to algorithmic complexity (Kolmogorov-Solomonoff-Chaitin complexity) [8, 9, 10, 11], denoted by $\mathbf{K}$, based on LZ compression scheme and asymptotically bounded by (Shannon) Entropy in a fashion that is similar to previous information and graph-theoretic approaches to chemical complexity [12, 13, 14, 15]. We conclude that Assembly Theory cannot offer a different or better explanation of selection or evolution than the connections already established [6].
2 Disassembling Assembly Theory: Main Concerns
2.1 The assembly Index is a compression algorithm of the LZ family
Despite the authors’ assertion in the Assembly Theory (AT) paper that AT and the assembly index are unrelated to algorithmic complexity, it is evident that AT is fundamentally encompassed within the realm of algorithmic complexity [16]. The assembly index, as proposed, seeks to gauge the complexity of an object based on the number of steps in its shortest copy-counting assembly pathway [2] via a procedure equivalent to LZ compression, which in turn is a computable approximation to $\mathbf{K}$.
At its core, the assembly index shares three key elements of entropy defined by Shannon himself [17], and shares the three main elements of the LZ family of compression algorithms [18, 19] that make LZ converge to entropy at the limit: the identification of repeated blocks, the usage of a dictionary containing repetitions and substitutions, and the ability to losslessly reconstruct the original object using its minimal LZ description. The parity in number of steps for compression and decompression using the assembly index effectively reduces the definition of Assembly Theory to the length of the shortest assembly pathway, which in turn is a loose upper bound of a resource-bounded approximation to algorithmic complexity $\mathbf{K}$ [16].
In contrast, more robust approximations to $\mathbf{K}$, capable of capturing blocks and other causal content within an object, have been proposed for purposes ranging from exploring cause-and-effect chains to quantifying object memory and characterising process content [20, 21]. These more advanced measures have found application in the same contexts and domains explored by Assembly Theory, encompassing tasks like distinguishing organic from non-organic molecules [15], investigating potential connections to selection and evolution [6, 7], the detection of bio and techno signatures [22, 23], and explorations into causality [24, 25]. Importantly, when applied to the data employed as evidence by AT, the more sophisticated measures consistently outperform the assembly index (see Suplementary Information and [16]). Thus, it becomes evident that Assembly Theory and its assembly index represent a considerably constrained version of compression algorithms, and a loose upper bound of $\mathbf{K}$.
This limitation is attributable to the authors’ exclusive consideration of computer programs adhering to the form of ‘Template Program A’, as follows:
Template Program A:
‘while end-of-object, do $N$ times print(repetitions) $+$ print (all remaining objects not found in repetitions)’
Any approximation to $\mathbf{K}$ that accounts for identical repetitions, including all known lossless statistical compression algorithms, can achieve equivalent or superior results, as demonstrated in the Suplementary Information and [16]. This alignment with ‘Template Program A’ effectively highlights the association of Assembly Theory with well-established principles of compression and coding theory, thereby refuting the initial claim of its authors to present a unique methodology. Moreover, the authors’ suggestion that their index may be generalised as a universally applicable algorithm for any object (including text) [26]further underscores the disconnection of Assembly Theory and their authors’ suggestion to reinvent text compression (e.g. ZIP or GZIP) from the current state of the art in the field of statistical and non-statistical compression beyond LZW.
It is worth noting that although there may be minor variations in the implementation details of the assembly index and the assembly number of which the authors themselves have proposed significantly different versions, we demonstrate that they are all qualitatively and quantitatively equivalent to the LZ compression algorithms introduced in the 1970s, such as LZ77/LZ78. See Sup. Inf. A.
In essence, the assembly index is fundamentally underpinned by the LZ encoding of the objects it measures. This reveals that the assembly index, and consequently AT, aligns more closely with $\mathbf{K}$ and the principles of traditional information theory (Shannon Entropy), and statistical data compression than the authors are willing to acknowledge [4].
The AT authors also assert that their index’s ability to differentiate between organic and non-organic compounds validates its natural applicability. However, when compared with other statistical indexes, including various compression algorithms, these alternative methods often engender similar or superior performances [16]. This undermines the claims made in favour of AT.
2.2 A compression algorithm is a mechanical procedure that corresponds to a physical process
The authors of AT argue that traditional data compression algorithms are an overly abstract process unsuited for modelling the construction (or assembly) of objects [26]. This view overlooks the practical and mechanistic nature of compression algorithms, particularly those in the LZ family. Since their introduction in 1977, LZ algorithms have been effectively used in detection, identification, clustering, and classification across various fields, including biology, chemistry, and medicine  [27, 15, 28], and to approximate algorithmic complexity [10, 29].
The argument presented by the proponents of Assembly Theory regarding the uncomputability of algorithmic complexity is deeply misguided. While it is true that $\mathbf{K}$ is semi-computable, computable algorithms like LZ77/LZ78/LZW have been widely used to approximate it. These algorithms have been applied in biology and chemistry, challenging the assertion that Assembly Theory represents a unique or superior approach  [15, 16]. Furthermore, in an earlier 1976 article [30], Lempel and Ziv defined an early version of their algorithm directly as a computable method to approximate algorithmic complexity, defining what is known as LZ complexity, which speaks to the initial motivation behind their now seminal compression algorithm.
The emphasis on the purported requirement of a Turing machine in the context of algorithmic complexity is a misdirected concern. A Turing machine is a synonym of ‘algorithm’ and an ‘algorithm’ is a synonym of a Turing machine, any rule or method that is algorithmic in nature, including all the computable methods in Assembly Theory, are algorithms, and, therefore, technically, Turing machines or programs that can run on a (universal) Turing machine. These facts are only a technicality unrelated to any putative advantage of Assembly Theory over other approaches to algorithmic complexity.
On the contrary, one of the distinctive features of algorithmic information theory (AIT), algorithmic complexity being one of its indexes, is that its importance and pervasiveness in mathematics, theoretical computer science, and complexity science are, in fact, established because of the invariance of its results with respect to the model of computation, whether abstract or physically implemented. More specifically, such statements from the authors of AT [26] indicate a fundamental misunderstanding of equivalence classes closed under reductions (in this case, computability classes), which are basic and pivotal concepts in computer science, complexity theory, and mathematics in general. It is evidence in support of this point that none of the formal proofs in the Sup. Inf. section A made any mention of any Turing machine.
In practice, computable approximations like LZ77/LZ78—which AT mimics without attribution—do not require a Turing machine, and have been widely and successfully used in clustering and categorisation  [27, 10], including in the successful separation of chemical compounds into organic and non-organic categories [15]. To say that a physical (computable/recursive) process like those affected by AT’s methods is not a compression algorithm because it does not resemble or correspond to the functioning of a Turing machine is as naively wrong and misplaced as saying a program written in Python cannot model the movement of a pendulum because nature does not run Python.
This type of argument also reflects a misunderstanding of the purposes of Turing machines and the foundations of computer science. As explained in a quote often attributed to Edsger W. Dijkstra,
“Computer science is no more about computers than astronomy is about telescopes.”
At its inception, Turing machines were defined within computer science as an abstraction of the concept of an algorithm, years before what we now know as computers were built.
In a published paper [4], the authors offered a proof of computability of their assembly index to distance themselves from Shannon Entropy and algorithmic (Kolmogorov) complexity. Their algorithm, categorised as Template Program A, is trivially computable and requires no proof of computability but also all other resource-bounded approximations to $\mathbf{K}$ are computable, including LZW that has been used for decades for this purpose [10]
2.3 Conflating object assembly process and causation directionality
The authors assume and present the sequential nature of their algorithm as an advantage [26]. Assembly Theory claims to be able to extract causal knowledge in the form of chains of cause and effect from its assembly pathways [26]. The authors’ central assumption is that each step of the basic Template Program A constitutes a cause-and-effect chain corresponding to how an object may have been physically assembled. This is no different from, and is indeed a restricted version of, a pushdown automaton capable of instantiating grammar compressor (which the authors rename as ‘assembly pathway’) and works exactly like an LZ algorithm.
The assumption of sequential assembly from instantaneous object snapshots is not based on any physical (biological, chemical, or other) evidence. In contrast, there is overwhelming evidence that this is not the case. For example, in the case of a genome sequence, all the regions of the genome sequence are exposed to selective forces simultaneously. Transcription factors, or genes that regulate other genes, do not get assembled or interact only sequentially. In chemistry, reactions do not happen on one side of a molecule first and then propagate to the other but happen in parallel. Reactions in time follow and are represented sequentially. However, throughout AT’s arguments, causal directionality in time is conflated with how an object may have been assembled from its instantaneous configuration.
This assumption of sequential assembly of an object based on causal direction is manifestly incorrect regarding how an object subject to selection assembles. In contrast, only by finding the generative mechanism of the object—such as the underlying set of (computable) mechanisms (of which the assembly index only takes into account one instance, the program that only counts copies)–and thus explaining the object in a non-trivial fashion, can one reproduce both the causal direction from the sequence of connected steps (see Fig. 1B) as well as how the object itself may have been assembled (which is not and cannot be in a sequential fashion).
Assembly Theory and its index cannot characterise this, for example, because the class of problems or functions recognised by pushdown automata (or generated by context-free grammars) is a proper subset of the class of recursive problems.
The authors of Assembly Theory have mistakenly claimed that algorithmic complexity and Turing machines are not, or cannot be, related to causality. This is incorrect [24, 31]. They were introduced as causal artifacts for studying mechanistic means for logical operations.
We have demonstrated that the assembly index requires a finite automaton to instantiate a restricted grammar (see Sup. Inf. A), a version of a specific-purpose machine, just as any other resource-bounded approximation to $\mathbf{K}$ would need to be instantiated (e.g. calculated, even by hand). In this case, the size of their formula or the size of the implementation of their computer program is the size of the special-purpose finite automaton, which is common and of fixed size for their calculations (which allows them to be discounted from the final length, just as it is from $\mathbf{K}$ or resource-bounded approximations).
The assembly index being reduced to a grammar compression scheme demonstrates not only that the assembly process’ complexity is dependent on the contingencies regarding which assembly spaces or paths were taken in the historical past stages to build the object, but also that this recursivity and dependence on past trajectories are subsumed into the concept of compressibility. Unlike algorithmic approaches, assembly theory does not incorporate any of the elements of causality such as perturbation and counterfactual analysis, as has been done in the context of algorithmic complexity, a generalisation that compared to AT reflects the simplistic nature of AT [25].
2.4 Lack of control experiments and absence of supporting empirical data beyond current domain knowledge
A critical examination of the AT methodology reveals significant shortcomings. Firstly, proponents of AT failed to conduct basic control experiments, a foundational aspect of introducing a new scientific metric. Benchmarking against established indices, particularly in coding and compression algorithms, is crucial to validating any new metric in the domain. Previous work on AT has never included meaningful experimental comparisons of the assembly index with other existing measures but schemes and mistaken theoretical arguments [4] (Figs. 1 and 2). Yet, we have shown that other algorithms, such as RLE, Huffman coding, and those based on entropy and dictionary-based methods, reproduced equivalent or superior results compared to the results published by the AT group [16].
Assembly Theory introduces a cutoff value in its index, purported to offer a unique perspective on molecular complexity, distinct from those based on algorithmic complexity. This value indicates when an object is more likely to be organic, alive, or a product of a living system. However, various indices and compression algorithms tested have yielded equivalent cutoff values (see [16]). In fact, such an Assembly Index threshold replicates previous results on molecular separation that employ algorithmic complexity. Thus, this overlap in results challenges the notion that Assembly Theory provides a unique tool for distinguishing between organic and non-organic matter. For instance, similar cutoff values have been derived in other studies, such as those mentioned in  [15], effectively separating organic from non-organic molecular compounds. While the authors of AT ignored decades of research in chemical complexity based on graph theory and the principles of Entropy [13, 12, 32, 14], all of which are equivalent to Assembly Theory, the depth of the purported experimental validation of AT was notably limited, especially when contrasted with more comprehensive studies. For instance, a more exhaustive and systematic approach, with proper control experiments involving over 15,000 chemical compounds [15] and employing algorithms from the LZ family (or Entropy) and others of greater sophistication (BDM), demonstrated the ability to separate organic from non-organic molecular compounds.
Turing’s motivation was to explore what could be automated in the causal mechanisation of operations (see Fig. 1B), for example, stepwise by hand using paper and pencil. The shortest among all the computer programs of this type is an upper bound approximation to $\mathbf{K}$. In other words, $\mathbf{K}$ cannot be longer than the length of this diagram. The concept of pathway complexity and copy number instantiated by the assembly index, is a restricted version of statistical nature of a Turing machine or finite automaton. Moreover, a Turing machine is a synonym of an algorithm and vice versa. It is a mistake to think of a Turing machine as a physical object or an object with particular properties or features (such as a head or tape). All the algorithms in assembly theory are Turing machines and viceversa.
Figure 1: A: The authors of AT have suggested that $\mathbf{K}$ would be proven to be contained in Assembly TheoryT [26]. This Venn diagram shows how Assembly Theory (AT) is connected and subsumed within algorithmic complexity or ($\mathbf{K}$) and within the group of statistical compression as proven in this paper (see Sup. Inf. A). B: Causal transition graph of a Turing machine with number 3019 (in Wolfram’s enumeration scheme [33]) with an empty initial condition found by using a computable method (e.g. CTM, [34]) to explain how the block-patterned string 111000111000 was assembled step-by-step based on the principles of $\mathbf{K}$ describing the state, memory, and output of the process as a fully causal mechanistic explanation. A Turing machine is simply a procedural algorithm and any algorithm can be represented by a Turing machine. By definition, this is a mechanistic process, and as physical as anything else, not an ‘abstract’ or ‘unrealisable’ process.
2.5 A circular argument cannot unify physics and biology
Central to Assembly Theory (AT) and the public claims made through their University Press Releases [5] is the connection that the authors of AT want to make to selection and evolution, maintaining that AT unifies physics and biology and explains and quantifies selection and evolution, both Darwinian and abiotic [2]. To demonstrate this connection, by assuming selectivity in the combination of linear strings $(P)$, the authors compare two schemes for combining linear strings $Q$, one random versus a non-random selection. This experiment yields observations of differences between the two $(R)$, and the authors conclude that selectivity $(S)$ exists in molecular assembly and, therefore, that AT can explain it.
Observing that a random selection differs from a non-random selection of strings, the authors use this as evidence for “the presence of selectivity in the combination process between the polymers existing in the assembly pool”.
Yet, assuming selectivity in the combination of strings and proceeding to say that a selection algorithm is expected to differ from a random one makes for a circular argument. Furthermore, the conclusion $(S)$ reaffirms the initial assumption $(P)$, resulting in a circular argument.
In a formal propositional chain, it can be represented as $P\implies Q\implies R\implies S\implies P$, where the conclusion merely restates the initial assumption, lacking any validation or verification, empirical or logical beyond a self-evident tautology.
Whether this sequential reasoning is relevant to how molecules are actually assembled is also unclear, given the overwhelming evidence that objects are not constructed sequentially and that object complexity in living systems is clearly not driven by identical copies only [35].
Nevertheless, the question of random versus non-random selection and evolution in the context of approximations to algorithmic complexity, including copy counting, was experimentally tested, empirically supported, and reported before in [6] following basic literature search and review, control experiments (comparison to other measures), and validation against existing knowledge in genetics and cell biology.
These circular arguments, lacking a foundation in chemistry, biology, or empirical evidence, lead the authors to propose the concept of “assembly time”. The suggestion is a time-scale separation between molecule production (assembly time) and discovery time, aiming to unify physics and biology. Time, however, has always been fundamental in evolutionary theory, and claims about the object history are the foundations of evolutionary theory. Assembly Theory revisits existing complexity science concepts described in the timeline provided in Fig. 2 but without acknowledgment or attribution.
Figure 2: A timeline of results in complexity science relevant to the claims and results from Assembly Theory. Assembly Theory renames several concepts such as dictionary trees as ‘assembly pathways’, heavily uses algorithmic probability in their reduction of combinatorial space arguments without attribution, and, as demonstrated, the assembly index and assembly number are equivalent to LZ lossless compression (proofs provided in the Sup. Inf. A)
3 Conclusions
The most popular examples used by the authors of Assembly Theory (AT) in their papers as illustrations of the way their algorithms work, such as ABRACADABRA and BANANA, are traditionally used to teach LZ77, LZ78, or LZW compression in Computer Science courses at university level. Dictionary trees have been used for pedagogic purposes in computer science for decades.
This would suggest some similarity between AT and compression algorithms, particularly those of a statistical nature based on recursion to previous states, copy counting, and the re-usage of patterns and repetitions.
In this paper, we have shown that the assembly index is a compression scheme belonging to the LZ family of compression algorithms, rather than only being similar. Diving into the details of their algorithms, we have revealed that they are equivalent to LZW and, therefore, to Shannon Entropy.
Our results demonstrate that an object with a low assembly index has high LZ compressibility (i.e., it is more compressible according to the LZ scheme) and, therefore, necessarily displays low entropy. In the opposite direction, an object with a high assembly index will have low LZ compressibility and, therefore, necessarily display high entropy. This means that Assembly Theory (AT), Shannon Entropy, and LZ compression algorithm are indistinguishable from each other, and the claim that those other approaches are incapable of dealing with their type of data, capture ‘structure’ or anything that could not have been already characterised by Shannon Entropy or LZW, as it has been done before, is inaccurate.
This paper also demonstrates that AT is subsumed into the theory and methods of Algorithmic Information Theory (AIT) as illustrated in Fig.1 1A and fully mathematically proven in the Sup. Inf. section A.
In addition, both theoretical concepts and empirical results in [1, 36, 4, 2] have been reported previously in earlier work in the context of AIT (from which algorithmic complexity is one of its indexes) for chemical processes in [15], and biology in [15, 37, 6, 38] but with comparison to other measures, including algorithms that go beyond statistical compression and connected to the concept of causal discovery.
Algorithmic complexity has been shown to be profoundly connected to causality [39, 40, 24, 41], not only for strings, but also it has been applied to images [38], networks [7, 42, 43, 44], vectors, matrices, and tensors in multiple dimensions [45, 46, 47]; and chemical structures [15].
In contrast to statistical compression schemes of which AT is a particular case, exploring other resource-bounded computable approximations to algorithmic complexity [21, 24], that consider aspects other than statistics, such as identical repetitions, may have more discriminatory power.
This is because they may tell apart cases in which a high assembly index value with an expected low copy number (or frequency of occurrence in the ensemble), low LZ compressibility and high entropy, would indicate that an object is more causally disconnected, independent, or statistically random, but actually is causally dependent due to being more compressible by exploiting, for example, short rewriting rules describing the object.
One positive aspect of Assembly Theory is that it offers a graph-like representation approach (albeit more convoluted than necessary) to approximate LZ compression in the specific context of molecular complexity, potentially making it more accessible to a broader and less technical community. However, it is inaccurate and a bad practice to ignore, omit or imply the absence of previous work (e.g. [15, 6, 48]), belittle and devalue the theories and ideas upon which their index is entirely dependable and connected with (Shannon Entropy, compression, and algorithmic Kolmogorov complexity), and make disproportionate public claims related to the contributions of AT (even calling it a theory that ‘unifies biology and physics’ and a ‘theory of everything’ by the media and the authors).
We agree with the authors on the importance of understanding the origin and mechanisms of the emergence of an open-ended generation of novelty [49, 41]. However, these efforts by the authors are undermined by hyperbolic claims.
We argue that the results and claims made by AT, in fact, highlight a working hypothesis that information and computation underpin the concepts necessary to explain the processes and building blocks of physical and living systems, concepts which the methods and frameworks of AT have been heavily inspired upon [50, 51].
4 Acknowledgements
Felipe S. Abrahão acknowledges the partial support of the São Paulo Research Foundation (FAPESP), Grant $2023$/$05593$-$1$.
    A.1 Mathematical framework
    As in [52], let $\left(\Gamma,\phi\right)$ denote either an assembly space or an assembly subspace. In the interests of of simplifying notation, one can refer to $\left(\Gamma,\phi\right)$ simply as $\Gamma$ where the edge-labellings given by $\phi$ are not relevant. From [52], we have it that $c_{\Gamma}\left(x\right)$ denotes the assembly index of the object $x$ in the assembly space $\Gamma$.
    Note that assembly spaces are finite. So, on the basis of our assumptions, we need to define a pathway assembly that can deal with arbitrarily large objects. To this end, let $\mathcal{S}=\left(\mathbf{\Gamma},\mathbf{\Phi},\mathcal{F}\right)$ be an infinite assembly space, where every assembly space $\Gamma\in\mathbf{\Gamma}$ is finite, $\mathbf{\Phi}$ is the set of the correspondent edge-labelling maps $\phi_{\Gamma}$ of each $\Gamma$, and $\mathcal{F}=\left(f_{1},\dots,f_{n},\dots\right)$ is the infinite sequence of embeddings (in which each embedding is also an assembly map as in [52]) that ends up generating $\mathcal{S}$. That is, each $\,f_{i}\colon\left\{\Gamma_{i}\right\}\subseteq\mathbf{\Gamma}\to\left\{\Gamma_{i+1}\right\}\subseteq\mathbf{\Gamma}\,$ is a particular type of assembly map that embeds a single assembly subspace into a larger assembly subspace so that the resulting sequence of nested assembly subspaces defines a total order $\preceq_{\mathcal{S}}$, where
[…]
    In other words, $\mathcal{S}$ is the nested family of all possible finite assembly spaces from the same basis set (i.e., the root vertex that represents the set of all basic building blocks) such that $\mathcal{S}$ is infinite and enumerable by a program. [^1] Notice that an alternative proof of the results presented in this paper can be achieved just with the (more general) assumption that the assembly space may be finite but only needs to be sufficiently large in comparison to the deceiving program in Lemma 1.
    Let $\gamma=z\dots y$ denote an arbitrary path from $z\in B_{\mathcal{S}}$ to some $y\in V\left(\mathcal{S}\right)$ in $\mathcal{S}$, where $B_{\mathcal{S}}$ is the basis set (i.e., the finite set of basic building blocks) of $\mathcal{S}$ and $V\left(\mathcal{S}\right)$ is the set of vertices of $\mathcal{S}$. Let $\gamma_{x}$ denote a rooted path from some $z\in B_{\mathcal{S}}$ to the object $x\in V\left(\mathcal{S}\right)$. All of these apply analogously to $\Gamma$ instead of $\mathcal{S}$.
    As in [52], let ${\Gamma^{*}}_{x}$ denote a minimum rooted assembly subspace of $\Gamma$ from which the assembly index $c_{\Gamma}\left(x\right)$ calculates the augmented cardinality (i.e., the number of vertices in the minimum rooted assembly subspace, except for those in the basis set $B_{\Gamma}$). The longest rooted paths $\gamma_{x}^{\max}$ of ${\Gamma^{*}}_{x}$ end in the arbitrary object/vertex $x\in V\left(\Gamma\right)$, i.e., $x$ is the terminal vertex of the path $\gamma_{x}^{\max}$ or, equivalently, $x$ is the target vertex of the latest oriented edge in $\gamma_{x}^{\max}$. The shortest rooted paths $\gamma_{x}^{\min}$ of ${\Gamma^{*}}_{x}$ end in the arbitrary object/vertex $x\in V\left(\Gamma\right)$, i.e., $x$ is the terminal vertex of the path $\gamma_{x}^{\min}$ or, equivalently, $x$ is the target vertex of the latest oriented edge in $\gamma_{x}^{\min}$.
    Let $\;A\#\colon\mathbf{X}\to\mathbb{N}\;$ be a function from the set $\mathbf{X}$ of possible ensembles (composed of objects) into the set of natural numbers. As in [2], this function defines the assembly number $A\#\left(X\right)$ of an ensemble $X$ such that
    $\begin{array}[]{lccc}A\#\colon&\mathbf{X}&\to&\mathbb{N}\\ &X&\mapsto&A\#\left(X\right)=A\left(N_{T},N,\left(n_{1},\dots,n_{N}\right),\left(a_{1},\dots,a_{N}\right)\right)\end{array}\text{ ,}$ (1)
    where:
    • $N_{T}$ is the total number of objects in the ensemble $X$;
    • $N$ is the total number of unique objects in the ensemble $N_{T}$;
    • each $n_{i}$ corresponds to the copy number of the (unique) object $x_{i}$ with $1\leq i\leq N$;
    • $a_{i}=c_{\Gamma}\left(x_{i}\right)$, where $\Gamma$ is the assembly space that contains the object $x_{i}$;
    • $A\left(\cdots\right)$ is the assembly of the ensemble $X$ calculated by
      $A\left(N_{T},N,\left(n_{1},\dots,n_{N}\right),\left(a_{1},\dots,a_{N}\right)\right)=\sum\limits_{i=1}^{N}e^{a_{i}}\left(\frac{n_{i}-1}{N_{T}}\right)\text{.}$
    According to the assumptions and claims in [1, 52, 2] (as also discussed in [16]), notice that the values of $N_{T}$, $N$, $n_{i}$, $a_{i}$ are retrievable by an effective procedure from a given (finite) ensemble $X$, computational methods which are proposed by Assembly Theory [1, 52, 2]. Thus, function $A\#$ is a recursive function (i.e., a function that can be calculated by an algorithm) given a finite ensemble $X$, which in turn the authors in [1, 52, 2] claim to be one of the advantages of Assembly Theory with respect to algorithmic complexity.
    Let $V\left(\Gamma\right)\right>$ be an encoding of the set $V\left(\Gamma\right)$ of vertices of the assembly space (or subspace) $\Gamma$. Notice that the encoding formalism and the ordering of the vertices belonging to $\Gamma$ are arbitrary. The only condition is that once the encoding method is (arbitrarily) chosen, it is fixed and remains the same for any possible assembly space (or subspace) $\Gamma$. Analogously, let $\gamma\right>$ denote an arbitrarily fixed encoding of a sequence/path $\gamma$ composed of contiguous oriented edges together with the starting vertex/object in the basis set $B_{\Gamma}\subseteq B_{\mathcal{S}}$, the sequence which defines the path $\gamma$ in $\Gamma$.
    We have it that the (prefix) algorithmic complexity, denoted by $\mathbf{K}\left(x\right)$, is the length of the shortest prefix-free (or self-delimiting) program that outputs the encoded object $x$ when this program is run according to a (universal) programming language.
    In accordance with the claims in [1, 36, 52, 2, 4], notice that the computability and feasibility of Assembly Theory’s methods directly imply the existence of algorithms that can enumerate the assembly spaces. Given any set of (physical, chemical, and/or biological) rules for assembling molecules that Assembly Theory arbitrarily chooses to be based on, one has it that for any particular pathway assembly resulting in a molecule, there is a correspondent algorithm for which one can apply the methods described in [1, 36, 52] to calculate the pathway probability, and so on. This existence is trivially guaranteed to hold as long as the calculation of the assembly index is recursive from a given assembly space, which is already an assumed condition in [1, 36, 52]; moreover, any upper bound for the computational resources necessary to compute the assembly index straightforwardly implies the existence of an upper bound for the computational resources necessary for a program to run. For every allowed generative process that can assemble objects and thereby build another object, there is an algorithm that performs this process. Conversely, for every object, one also has it that there is a corresponding generative process allowed by Assembly Theory, a process which is effected by an algorithm. Therefore, according to the assumptions in [1, 36, 52, 2, 4], there is a formal theory $\mathbf{F}$ that contains Assembly Theory and all the procedures effected by algorithms, including the chosen method for calculating the assembly index of an object, the program that decides whether or not the criteria for building the assembly spaces are met, and so on.
    A.2 Assembly theory is an approximation to algorithmic complexity
    One can demonstrate in Lemma 1 and Corollary 2 that the assembly index $c_{\Gamma}\left(y\right)$ (or the assembly number $A\#\left(X\right)$ ) is an approximation to algorithmic complexity $\mathbf{K}\left(y\right)$ (or, respectively, $\mathbf{K}\left(X\right)$) that in general overestimates its value as much as one wishes.
    As in [16], the main avenue to achieving the following theoretical results is to construct a randomly generated program that receives a formal theory (which contains all the procedures and statistical criteria in Assembly Theory) as input. Then, it searches for an object $y$ in an assembly space with a sufficiently high assembly index $c_{\Gamma}\left(y\right)$ so as to overcome any arbitrarily chosen error margin $k$ between $c_{\Gamma}\left(y\right)$ and the algorithmic complexity $\mathbf{K}\left(y\right)$.
    Lemma 1.
    Let $\mathcal{S}$ be enumerable by an algorithm. Let $\mathbf{F}$ be an arbitrary formal theory that contains Assembly Theory, including all the decidable procedures of the chosen method for calculating the assembly index of an object for a nested subspace of $\mathcal{S}$, and the program that decides whether or not the criteria for building the assembly spaces are met. Let $k\in\mathbb{N}$ be an arbitrarily large natural number. Then, there are $\Gamma\subset\mathcal{S}$ and $y\in V\left(\Gamma\right)$ such that
    $\mathbf{K}\left(y\right)+k\leq c_{\Gamma}\left(y\right)\text{ ,}$ (2)
    where the function $\;c_{\Gamma}\colon\Gamma\subset\mathcal{S}\to\mathbb{N}\;$ gives the assembly index of the object $y$ in the assembly space $\Gamma$ (or $\mathcal{S}$).
    Proof.
    Let $p$ be a bit string that represents an algorithm that receives $\mathbf{F}$ and $k$ as inputs. Then, it calculates $\left|p\right|+\left|\mathbf{F}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+k$ and enumerates $\mathcal{S}$ while calculating $c_{\Gamma}\left(x\right)$ of the object (or vertex) $x\in V\left(\Gamma\right)\subset V\left(\mathcal{S}\right)$ at each step of this enumeration. Finally, the algorithm returns the first object $y\in V\left(\mathcal{S}\right)$ for which
    $\left|p\right|+\left|\mathbf{F}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+k+\mathbf{O}(1)\leq c_{\Gamma}\left(y\right)$ (3)
    holds. In order to demonstrate that $p$ always halts, just note that $\mathcal{S}$ is infinite computably enumerable. Also, for any value of $c_{\Gamma^{\prime}}\left(z\right)$ for some $z\in V\left(\Gamma^{\prime}\right)\subset V\left(\mathcal{S}\right)$, there is only a finite number of minimum rooted assembly subspaces (starting on any object in $B_{\mathcal{S}}$ and ending on $z$) whose augmented cardinality is $c_{\Gamma^{\prime}}\left(z\right)$, where $B_{\mathcal{S}}$ is the basis (i.e., the finite set of basic building blocks [1]) of $\mathcal{S}$. This implies that there is an infinite number of distinct values of $c_{\Gamma^{\prime}}\left(z\right)$. Now, let $\mathrm{p}_{y}=\left<k,\mathbf{F},p\right>$. Finally, from Equation (3) and basic properties in AIT, we have it that
    $\mathbf{K}\left(y\right)+k\leq\left|\mathrm{p}_{y}\right|+k+\mathbf{O}(1)\leq\left|p\right|+\left|\mathbf{F}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+k+\mathbf{O}(1)\leq c_{\Gamma}\left(y\right)$ (4)
    holds for some sufficiently large $k$. ∎
    From this proof of Lemma 1, one can directly conclude that:
    Corollary 2.
    There are infinitely many $\Gamma\subset\mathcal{S}$ for which Equation (2) in Lemma 1 holds.
    Proof.
    From the proof of Lemma 1 we know that there are infinitely many distinct values of $c_{\Gamma^{\prime}}\left(z\right)$. Let $p_{2}$ be a slight variation of program $p$ in the proof of Lemma 1 such that, in addition to $\mathbf{F}$ and $k$, it also receives a natural number $i\in\mathbb{N}$ as input. Then, it enumerates $\mathcal{S}$ while calculating $c_{\Gamma}\left(x\right)$ of the object (or vertex) $x\in V\left(\Gamma\right)\subset V\left(\mathcal{S}\right)$ at each step of this enumeration. In the last step, program $p_{2}$ returns the $i$-th first object $y\in V(\mathcal{S})$ for which
    $\left|p\right|+\left|\mathbf{F}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+k+\mathbf{O}\left(\log_{2}\left(i\right)\right)+i+\mathbf{O}(1)\leq c_{\Gamma}\left(y\right)$ (5)
    holds. (Notice that Equation (5) differs from Equation (3)). Then, for each $i>0$, one will obtain a distinct object $y$ such that
    $\mathbf{K}\left(y\right)+k\leq\left|p\right|+\left|\mathbf{F}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+k+\mathbf{O}\left(\log_{2}\left(i\right)\right)+i+\mathbf{O}(1)\leq c_{\Gamma}\left(y\right)$ (6)
    holds. ∎
    A.2.1 Assembly number and algorithmic complexity
    Now we can extend Lemma 1 to the assembly number $A\#$, a value which is defined upon an ensemble $X$. The main idea of Lemma 3 is that the class of all possible ensembles of objects that assembly spaces can possibly cover is also enumerable by an algorithm, based on the fact that assembly spaces are enumerable by an algorithm. In this manner, as done in Lemma 1, one can make the error margins (in comparison to algorithmic complexity) as large as one wishes.
    Lemma 3.
    Let $\mathcal{S}$ be enumerable by an algorithm. Let $\mathbf{F}$ be an arbitrary formal theory that contains Assembly Theory, including: all the decidable procedures of the chosen method for calculating the assembly index $c_{\Gamma}$ of an object for a nested subspace of $\mathcal{S}$ and the assembly number $A\#$ of a given ensemble $X$; the program that decides whether or not the criteria for building the assembly spaces are met; and the program that decides whether or not the criteria for building ensembles of objects are met. Let $k\in\mathbb{N}$ be an arbitrarily large natural number. Then, there is $y\in\mathbf{X}$ such that
    $\mathbf{K}\left(X\right)+k\leq A\#\left(X\right)\text{ ,}$ (7)
    where the function $\;A\#\colon\mathbf{X}\to\mathbb{N}\;$ gives the assembly number of the ensemble $X\in\mathbf{X}$.
    Proof.
    Remember the definition of assembly number $A\#$ such that $A\#\left(X\right)=\sum\limits_{i=1}^{N}e^{a_{i}}\left(\frac{n_{i}-1}{N_{T}}\right)$, where the parameters $N_{T}$, $N$, $n_{i}$, $a_{i}$ are computably retrievable from a given ensemble $X$. Since $\mathcal{S}$ is infinite computably enumerable and the theory $F$ can always decide whether or not an arbitrary set $Y$ is an ensemble from which $A\#$ can be calculated, one has it that the set $\mathbf{X}$ (of all possible finite ensembles) is infinite computably enumerable. From here on, the proof follows analogously to the proof of Lemma 1. ∎
    In the same manner as in Corollary 2 with respect to Lemma 1, one can obtain the following corollary:
    Corollary 4.
    There are infinitely many $X\in\mathbf{X}$ for which Equation (7) in Lemma 3 holds.
    Proof.
    Analogous to the proof of Corollary 2. ∎
    The reader is invited to note that the latter Lemma 3 and Corollary 4 can also be generalised for any recursive function $f$ defined on top of the assembly index and on an arbitrary observable (for which Assembly Theory, in particular, would include a formal criterion to decide whether or not function $f$ can be calculated from this observable).
    A.3 Assembly theory is an encoding/compression scheme
    A.3.1 Assembly index defines an encoding scheme
    The following results show that the assembly index $c_{\Gamma}\left(y\right)$ is a parameter that can be employed to encode an object $y$. Theorem 5 demonstrates that any arbitrarily chosen encoding of the set of vertices (in any arbitrary ordering) of a minimum rooted assembly space is sufficient to encode the highest-assembly-index object. In other words, the sheer collection of objects in a minimum rooted assembly subspace carries enough information to allow one to always pick the object with the highest assembly index value, even though the ordering of the objects in this collection can be completely arbitrary in the first place. This demonstrates that Assembly Theory is one of the possible methods to encode objects.
    Theorem 5.
    Let $\mathcal{S}$ be enumerable by an algorithm. Let $\mathbf{F}$ be an arbitrary formal theory that contains Assembly Theory, including all the decidable procedures of the chosen method for calculating the assembly index of an object for a nested subspace of $\mathcal{S}$, and the program that decides whether or not the criteria for building the assembly spaces are met. Let $\Gamma\subset\mathcal{S}$ with $y\in V\left(\Gamma\right)$ be arbitrary. Then, one has it that $V\left({\Gamma^{*}}_{y}\right)\right>$ encodes the object $y$ such that
    $\mathbf{K}\left(y\right)\leq\mathbf{K}\left(\left<V\left({\Gamma^{*}}_{y}\right)\right>\right)+\mathbf{O}(1)$ (8)
    holds, where: $\;c_{\Gamma}\colon\Gamma\subset\mathcal{S}\to\mathbb{N}\;$ gives the assembly index of the object $y$ in the assembly space $\Gamma$ (or $\mathcal{S}$); ${\Gamma^{*}}_{y}$ is the minimum rooted assembly subspace of $\Gamma$ from which the assembly index $c_{\Gamma}\left(y\right)$ is calculated; and $V\right>$ is an arbitrarily fixed encoding of a set $V$ of vertices.
    Proof.
    Remember that ${\Gamma^{*}}_{y}$ denotes the minimum rooted assembly subspace of $\Gamma$ from which the assembly index calculates the augmented cardinality that defines the value $c_{\Gamma}\left(y\right)$. Notice that since ${\Gamma^{*}}_{y}$ is minimal, then there is a simple algorithm that can always calculate the integer value of $c_{\Gamma}\left(y\right)$ given $\mathbf{K}\left(\left<V\left({\Gamma^{*}}_{y}\right)\right>\right)$ as input. Remember that, from our definitions and conditions in Section A.1, $V\left({\Gamma^{*}}_{y}\right)\right>$ denotes an encoding of the set of vertices of ${\Gamma^{*}}_{y}$, an encoding method which is arbitrary, previously chosen, and does not depend on the choice of the assembly (sub)space or object. Now, one can exploit the computability of Assembly Theory and construct a program $p_{3}$ that receives $\mathbf{F}$ and $V\left({\Gamma^{*}}_{y}\right)\right>$ as inputs. By using $\mathbf{F}$ to decide whether or not an arbitrary $\Gamma^{\prime}$ satisfies the criteria of being an assembly subspace, program $p_{3}$ enumerates the finite set $\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$ of all possible assembly subspaces whose set of vertices is exactly $V\left({\Gamma^{*}}_{y}\right)$. Next, using $\mathbf{F}$ to decide whether or not an arbitrary $\Gamma^{\prime}$ satisfies the criteria of being a minimum rooted assembly subspace whose longest path ends in the same last vertex of $\Gamma^{\prime}$, program $p_{3}$ searches in $\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$ for the first assembly subspace $\Gamma^{\prime\prime}$ that is minimal. Finally, it returns the object $y^{\prime}\in V\left(\Gamma^{\prime\prime}\right)$ at which the longest rooted paths in $\Gamma^{\prime\prime}$ end (i.e., $y^{\prime}$ is the terminal vertex of these longest rooted paths). In order to achieve the desired proof, one has yet to show that the object built at the end of the longest paths in the minimum rooted assembly subspaces is unique and $y^{\prime}=y$. To this end, one can employ the minimality of ${\Gamma^{*}}_{y}$. First, suppose there is a minimum $\Gamma_{2}\in\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$ that ends at the object $z$ such that $\Gamma_{2}\neq\Gamma^{\prime\prime}$ and $z\neq y^{\prime}$. Then, by construction of the program $p_{3}$, one would have it that $y^{\prime}$ belongs to at least one of the rooted paths in $\Gamma_{2}$. Hence, there would be a rooted path that leads to $y^{\prime}$ that is shorter than the one that leads to $z$ because by construction one has it that the path that leads to $z$ is the longest one; and if there were another terminal vertex $w$ in $\Gamma_{2}$ whose rooted path has the same length of that of $z$, then $\Gamma_{2}$ would not be a minimum subspace. As a consequence, there would be a rooted assembly subspace ${\Gamma_{2}^{*}}_{y^{\prime}}\subseteq\Gamma_{2}$ whose augmented cardinality is smaller than $c_{\Gamma^{\prime\prime}}\left(y^{\prime}\right)$. However, by construction, we have it that $c_{\Gamma^{\prime\prime}}\left(y^{\prime}\right)$ is minimal, which leads to a contradiction. Therefore, since by hypothesis we have it that ${\Gamma^{*}}_{y}$ is minimal and ${\Gamma^{*}}_{y}\in\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$, we conclude that either $\Gamma_{2}=\Gamma^{\prime\prime}={\Gamma^{*}}_{y}$ or $z=y^{\prime}=y$, and hence that, in any case, program $p_{3}$ produces $y$ as output. Finally, the proof of Equation (8) follows from basic inequalities in AIT. ∎
    Notice that $V\left({\Gamma^{*}}_{y}\right)\right>$ only encodes the membership relation for the set, and does not need to carry any information about its internal structure that results in the assembly space from which it only takes the set of vertices. In fact, a straightforward analogous proof to the one in Theorem 5 can be applied to demonstrate that not only is the final object encodable by $V\left({\Gamma^{*}}_{y}\right)\right>$, but also the longest paths in the respective minimum rooted assembly subspace. A sequence of edges in each of these longest paths constitutes a way to parse the object (e.g., a string), i.e., to decompose it into contiguous subsequences. This idea will become clearer in the next Sections A.3.2 and A.3.3. It also constitutes the compression scheme underlying LZ algorithms.
    Nevertheless, one can now demonstrate that the opposite also holds, that is, from encoding one of the longest paths in a minimum rooted assembly subspace, one can always directly retrieve the minimum rooted assembly subspace, and hence the terminal vertex that corresponds to the object to be assembled:
    Corollary 6.
    Let $\mathcal{S}$ be enumerable by an algorithm. Let $\mathbf{F}$ be an arbitrary formal theory that contains Assembly Theory, including all the decidable procedures of the chosen method for calculating the assembly index of an object for a nested subspace of $\mathcal{S}$, and the program that decides whether or not the criteria for building the assembly spaces are met. Let $\Gamma\subset\mathcal{S}$ with $y\in V\left(\Gamma\right)$ and $\gamma_{y}^{\max}$ be arbitrary, where $\gamma_{y}^{\max}$ is one of the longest rooted paths in ${\Gamma^{*}}_{y}$ that has $y$ as its terminal vertex and ${\Gamma^{*}}_{y}$ is the minimum rooted assembly subspace of $\Gamma$ from which the assembly index $c_{\Gamma}\left(y\right)$ is calculated. Then, one has it that $\gamma_{y}^{\max}\right>$ encodes the object $y$ such that
    $\mathbf{K}\left(y\right)\leq\mathbf{K}\left(\left<\gamma_{y}^{\max}\right>\right)+\mathbf{O}(1)$ (9)
    holds, where $\gamma\right>$ is any arbitrarily fixed encoding of a sequence $\gamma$ composed by contiguous oriented edges together with the starting vertex/object in the basis set $B_{\Gamma}\subseteq B_{\mathcal{S}}$, the sequence which defines the path $\gamma$.
    Proof.
    Let $p_{3}$ be the program used in the proof of Theorem 5. Let $p_{4}$ be a program that receives $\mathbf{F}$, program $p_{3}$, and $\gamma_{y}^{\max}\right>$ as inputs. Then, by using $\mathbf{F}$ to decide whether or not an arbitrary $\Gamma^{\prime}$ satisfies the criteria of being a minimum rooted assembly subspace, enumerates all possible minimum rooted assembly subspaces that have $\gamma_{y}^{\max}$ as one of its longest rooted paths. In the next step, program $p_{4}$ picks the first of these possible minimum rooted assembly subspaces and encodes it in the form $V\left({\Gamma^{*}}_{y}\right)\right>$. Finally, program $p_{4}$ returns the output of program $p_{3}$ with $\mathbf{F}$ and $V\left({\Gamma^{*}}_{y}\right)\right>$ as inputs. The proof of Equation (9) follows from Equation (8) and basic inequalities in AIT. ∎
    A.3.2 Assembly index defines a grammar of compression that is bounded by LZ compression and can at most approximate the Shannon Entropy rate
    In this section, we investigate how well the assembly index can also produce an encoding whose size should be as small as possible in comparison to the size of the original object. In other words, we investigate how the assembly index performs as a compression method. Indeed, we show that Assembly Theory is not only an encoding but also a compression scheme, in particular one that decomposes (i.e., parses) the object into contiguous building blocks, as is done by other methods like LZ compression algorithms and compressing grammars.
    As presented in Section A.1, recall that the assembly index $c_{\Gamma}\left(x\right)$ of an object $x$ is defined as the size of the smallest acyclic quiver (directed multigraph) with edge-labellings from the very set of vertices of this multigraph, a quiver which is called the assembly (sub)space, denoted by $\left(\Gamma,\phi\right)$, or just $\Gamma$ where the edge-labellings given by $\phi$ are not relevant. The “tree”-like structure of the assembly space maps the generation of an object from a set of “basic objects” in $B_{\mathcal{S}}$, which is specified to be finite. Each vertex represents an object, and the directed edge relation specifies the generation hierarchy (i.e., the label of the oriented edge) of the object to be combined with the previous object (i.e., the source vertex to which the respective outgoing labelled edges are connected) in order to assemble a new object (i.e., the target vertex to which the respective incoming labelled edges are connected).
    Now, given that these minimum rooted assembly subspaces are finite and acyclic, the number of objects they can encode is finite. Therefore, there is a CFG (Context-Free Grammar) that generates all these objects from an alphabet of terminal symbols, which correspond to the set of basic objects. Grammar-based compression algorithms have been an active area of research [53, 54, 55, 56]. The problem of finding the smallest grammar has been studied for several decades and is known to be NP-Complete [57, 54]. Furthermore, the size of these compressing grammars is lower bounded by LZ encoding [55] (disregarding time complexity), which means that they converge to LZ in the optimal case for ergodic and stationary stochastic processes, and therefore are bounded by the noiseless compression limit [58].
    It is straightforward to show the existence of such a CFG that encodes all the objects generated by an assembly space. In the following, we will investigate it further to demonstrate that for any minimum $\Gamma$ that generates a string $x$, there exists a grammar that encodes this $\Gamma$. In other words, the grammar generates $x$ and manifests a correspondence between: the basic objects of both the grammar (terminal characters) and the assembly subspace; non-basic objects and vertices of the quiver with the sets of non-terminal characters; and edges of the quiver with derivation rules of the grammar.
    Theorem 7.
    Let $\Gamma=\langle V_{\Gamma},E_{\Gamma},s_{\gamma},t_{\gamma}\rangle$ be the acyclic finite quiver with basic objects $\Sigma_{G}\subset V_{\Gamma}$ such that was defined to compute the assembly index $c_{\Gamma}(x)$. There exists a CFG $G=\langle V_{G},\Sigma_{G},R,S\rangle$ such that $V_{\Gamma}-\Sigma=V_{G}$, $\Sigma_{\Gamma}=\Sigma_{G}$ and $|E_{\Gamma}|=|R|$.
    Proof.
    Let $\Gamma$ be the acyclic finite quiver that encodes the generation of the string $x$. First, we define the set of terminal symbols in $G$ as the set of basic objects in $\Gamma$, $\Sigma_{G}=\Sigma_{\Gamma}=\Sigma$. Let $V_{G}=(V_{\Gamma}-\Sigma)\cup\{S\}$ be the set of non-terminal symbols.
    Now, we construct $R$ in the following way: We start with $S\rightarrow x$. Now, for each $e_{i}\in E$ such that $s_{\gamma}(e_{i})=x_{i}$ and $s_{\gamma}(e_{i})=y_{i}$, we define the rule as $y_{i}\rightarrow x_{i}z_{i}$ or $y_{i}\rightarrow x_{i}z_{i}$, $z_{i}\in V_{\Gamma}$, whichever concatenation of combinations builds the object encoded by vertex $y_{i}$. This is possible since each “assembly step” can only combine two objects. If the combination was unrestricted, then the assembly index of strings would always be one, as you can combine all the characters in the first step. For any other limit we can use a rule with more symbols.
    It is easy to prove by induction over the number of vertices $|V_{\Gamma}|$ that $G$ produces $x$. If $|V_{\Gamma}|=1$, then $x$ must be basic in $\Gamma$, therefore terminal in $G$. It follows that $S\rightarrow x$ generates $x$. Now we assume that the theorem holds for $n$ or less; consider the quiver $V_{\Gamma}={x_{1},\dots,x_{n},x}$ so that $\Gamma$ generates $x$. By hypothesis there exists an edge that, for some $x_{i}$, goes to $x$. Therefore there is the rule $x\rightarrow x_{i}z_{i}$ or $X\rightarrow z_{i}x_{i}$ such that $x$ is either concatenation. By hypothesis, the subgrammars defined by removing $x$ and adding the rule $S\rightarrow z_{i}$ and $S\rightarrow x_{i}$ must generate $z_{i}$ and $x_{i}$. Therefore we have the rules to generate $z_{i}$ and $x_{i}$; thus with the rule $x\rightarrow x_{i}z_{i}$ or $x\rightarrow z_{i}x_{i}$ we can generate $x$. Then, since $S\rightarrow X$, $\Gamma$ generates $x$.
    Finally, we remove the rule $S\rightarrow x$ and replace $x$ by $S$, which is equivalent to just defining the starting symbol as being $x$. By construction, the stated equities between the sets hold.
    The grammar defined in Theorem 7 demonstrates that any representation of a string’s assembly by an acyclic finite directed graph can be encoded into a Context-Free Grammar (CFG) in Chomsky Normal Form (CNF), capturing the same information regarding the relationship of how a string is constructed from basic components. The assembly index is defined as the number of non-basic vertices in the graph, while the size of the grammar is defined as the number of rules, which is equivalent to the number of nonterminals if there are no useless rules [55]. Let’s denote this size by $|G|$. Cronin et al. define the assembly space as a quiver. It is not clear that there is any necessity for having two or more edges linking the same pair of vertices in an assembly space, but for completeness’ sake, we will denote by $G^{*}$ the grammar with any useless rules removed. We will show that there is a simple linear relationship between the assembly index $c_{\Gamma}(x)$ and the size of the grammar $|G^{*}|$.
    Corollary 8.
    Let $x$ be a string with minimal assembly space $\Gamma$ and $G$ be the CFG in Theorem 7. Then
    • $c_{\Gamma}(x)=|G^{*}|$.
[…]
    Proof.
    The first equation is by construction of $G$. The second is a direct consequence of the first and [55, Theorem $1$]. ∎
    With the stated corollary we have proven that the assembly index for strings is equivalent to the compression length of a compressing grammar and, in the optimal case it can only converge towards LZ compression [30, 59, 60] and thus is bounded by Shannon Entropy by the (noiseless) Source Coding Theorem [61, 58]. It is important to note that the inequality arises when the grammar, thus the pathway assembly, is not optimal at identifying the ”building blocks” of $x$ given the same alphabet of basic objects. For compressing grammars in particular, several implementation-specific entropy bounds were found in [56].
    Now, for higher-dimensional objects, if the construction is recursive, it is possible to find a one-dimensional rule that encodes the object’s assembly. For example, for a two-dimensional object, we can construct a grid of fixed size around the object and incorporate the positional information into the grammar’s rules while maintaining a linear relationship between the assembly index and the size of the grammar. The rule $X\rightarrow ApB$ can define how object $B$ is added to object $A$ at position $p$, where $p$ corresponds to a square in the grid, to generate object $X$.
    A.3.3 Assembly index is a version of LZ compression
    In Section A.3, we have demonstrated that either the set of vertices of the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ or at least one of the longest rooted paths in this ${\Gamma^{*}}_{y}$ is sufficient to encode the object $y$. These results employ general algorithmic methods so that they are independent of the choice of the programming language and encoding-decoding schemes. In Theorem 7, we have demonstrated that every finite assembly space can be generated by a corresponding context-free grammar (CFG), which reveals a way one can compress an assembly space (e.g., the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ that generates the object $y$) into the minimum set of rules of this CFG. These results together reveal the existence of some LZ-based method to encode the object and its assembling process.
    Indeed, the proof of the following Theorem 9 constructs such a LZ compression scheme. In practice, the major characteristic that distinguishes assembly index compression from other traditional LZ schemes, is that: in the latter case, one would parse the string into contiguous smaller strings (or codewords) whose resulting unidimensional concatenation directly represents the string; while in the former case, one would parse the object into a non-unidimensional structure where each block is “concatenated” according to a directed acyclic multigraph. Thus, as shown in Theorem 9, a LZ compression scheme from the assembly index requires a way to compress this non-linear structure (in particular, a directed acyclic multigraph) back into a linear sequence of codewords.
    For LZ-based schemes, one has a sequence of encoded tuples of finite length. Typically, each one of these tuples contains at least one pointer (e.g., in the case of LZ78787878 and LZW, called indexes) that represents to another tuple. Also, it contains information about some elements of the basic alphabet (e.g., in the case of LZ78787878 and LZW, it contains the codeword for only one symbol to be concatenated at the end of the element the pointer is referring to). In the same fashion, one expects that the sequence of edges that constitutes an assembling path (from the basis set $B_{\Gamma}$ to the terminal object/vertex in a minimum rooted assembly subspace ${\Gamma^{*}}_{y}$) can be converted into a sequence of respective encoded tuples that contains pointers, codewords for other objects, etc. One can think of these assembling paths being, for example, the sequence defined by either a longest rooted path $\gamma_{x}^{\max}$ (as investigated in Section A.3.1) or a shortest rooted path $\gamma_{x}^{\min}$, where both belong to the minimum rooted assembly subspace ${\Gamma^{*}}_{x}$.
    The pointers in LZ compression schemes traditionally refer to encoded tuples that have occurred before. For Assembly Theory, the pointers may refer to other tuples that have not occurred before in an assembling path, but are part of the minimum rooted assembly subspace ${\Gamma^{*}}_{x}$, that is, part of a “tree”-like structure. We demonstrate in Theorem 9 that the minimality of ${\Gamma^{*}}_{x}$ in fact guarantees that such a scheme, where pointers may not refer to the unidimensional past subsequences, allows unique decoding, despite the non-unidimensional structure of assembly spaces.
    Theorem 9 (LZ encoding from the assembly index).
    Let $\mathcal{S}$ be enumerable by an algorithm. Let $\mathbf{F}$ be an arbitrary formal theory that contains Assembly Theory, including all the decidable procedures of the chosen method for calculating the assembly index of an object for a nested subspace of $\mathcal{S}$, and the program that decides whether or not the criteria for building the assembly spaces are met. Let $\Gamma\subset\mathcal{S}$ with $y\in V\left(\Gamma\right)$, $\gamma_{y}^{\min}$, and $\gamma_{y}^{\max}$ be arbitrary. Then, one has it that the sequence $S_{AT}\left(y\right)$ composed of $\left|\gamma_{y}^{\min}\right|$ encoded tuples in the form $c,i_{t},v_{l}\right>$ (except for the 00-th initial tuples which are $t\right>$, $\left|\gamma_{y}^{\max}\right|\right>$, $B_{\Gamma}\right>$, and $v_{B}\right>$, respectively, where $v_{B}\in B_{\Gamma}$ and $t\in\mathbb{P}\left(\mathcal{T}\right)$) encodes the object $y$ such that
    $\mathbf{O}\left(\left|\gamma_{y}^{\min}\right|\right)\leq$ (10)
    $\left|S_{AT}\left(y\right)\right|\leq$
    $\mathbf{O}\left(2^{3\,\left|\gamma_{y}^{\max}\right|+6}+\left|B_{\Gamma}\right|\log\left(\left|B_{\Gamma}\right|\right)\right)+\mathbf{O}\left(\log\left(\left|\gamma_{y}^{\max}\right|\right)\right)+$
    $\mathbf{O}\left(\left(\left|B_{\Gamma}\right|+1\right)\,\log\left(\left|B_{\Gamma}\right|\right)\right)+\left|\gamma_{y}^{\min}\right|\Big{(}\mathbf{O}\left(3\,\left|\gamma_{y}^{\max}\right|+6\right)+\log\left(\left|\gamma_{y}^{\min}\right|\right)\Big{)}\leq$
    $\mathbf{O}\left(2^{3\,c_{\Gamma}\left(y\right)+6}+\left|B_{\Gamma}\right|\log\left(\left|B_{\Gamma}\right|\right)\right)+\mathbf{O}\left(\log\left(c_{\Gamma}\left(y\right)\right)\right)+$
    $\mathbf{O}\left(\left(\left|B_{\Gamma}\right|+1\right)\,\log\left(\left|B_{\Gamma}\right|\right)\right)+c_{\Gamma}\left(y\right)\Big{(}\mathbf{O}\left(3\,c_{\Gamma}\left(y\right)+6\right)+\log\left(c_{\Gamma}\left(y\right)\right)\Big{)}$
    holds, where:
    • $\left|\gamma_{y}^{\min}\right|$ is the length of the rooted path $\gamma_{y}^{\min}$, which is one of the shortest rooted paths with $y$ as terminal vertex that belongs to the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$;
    • $\left|\gamma_{y}^{\max}\right|$ is the length of the rooted path $\gamma_{y}^{\max}$, which is one of the longest rooted paths with $y$ as terminal vertex that belongs to the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$;
    • $\left|\gamma_{y}^{\max}\right|\right>$ is an encoding of the number $\left|\gamma_{y}^{\max}\right|$;
    • $B_{\Gamma}\right>$ is an encoding of the basis set $B_{\Gamma}$;
    • $v_{B}\right>$ is an encoding of the vertex $v\in B_{\Gamma}$ which is the initial object/vertex of $\gamma_{y}^{\min}$;
    • $c$ is the counter of the tuple in the sequence $S_{AT}\left(y\right)$ (i.e., the natural number that indicates the position of the tuple) with $1\leq c\leq\left|\gamma_{y}^{\min}\right|$;
    • $t\right>$ is an encoding of $t\in\mathbb{P}\left(\mathcal{T}\right)$, where $\mathbb{P}\left(\mathcal{T}\right)$ is the powerset of possibles edges of the minimum directed acyclic multigraph $\mathcal{T}$ into which any possible minimum rooted assembly subspace whose longest rooted path has length $\left|\gamma_{y}^{\max}\right|$ can be embedded, and all the most distant vertices from the source can only be the vertices in $B_{\Gamma}$.
    • $i_{t}$ is the index/pointer that refers to the corresponding edge in $\mathcal{T}$;
    • $v_{l}$ is another index/pointer that refers to the edge $i_{t}$ whose source vertex $s\left(i_{t}\right)=v\in V\left(\Gamma\right)$ is the vertex that is the label of the edge $i_{t}$ when $\Gamma$ is embedded into $\mathcal{T}$;
    • $\left|S_{AT}\left(y\right)\right|$ denotes the size in bits of the encoded sequence $S_{AT}\left(y\right)$.
    • $\;c_{\Gamma}\colon\Gamma\subset\mathcal{S}\to\mathbb{N}\;$ gives the assembly index of the object $y$ in the assembly space $\Gamma$ (or $\mathcal{S}$);
    The encoding-decoding scheme for Theorem 9.
    Let $p_{5}$ be a program that receives a sequence $S_{AT}\left(y\right)$ in the form
[…]
    as input. Then, it constructs the directed acyclic multigraph (DAmG) $\mathcal{T}$ of height $\left|\gamma_{y}^{\max}\right|$, and enumerates all the possible minimum rooted assembly subspaces contained in $\mathcal{T}$: starting from those that are isomorphic to $\gamma_{y}^{\max}$ with any possible combination of initial vertex/object in $B_{\Gamma}$; and then in a dovetailing procedure, indexing first those subspaces that differ (from $\gamma_{y}^{\max}$) by the least amount of additional edges. In the next step, program $p_{5}$ reads the rest of the sequence $S_{AT}\left(y\right)$, extracting from the encoded tuples the set [^2] Note that some of these members may actually refer to the same index. Thus, the size of this set may be much smaller than $2\left|\gamma_{y}^{\min}\right|$. $I=\left\{{i_{t}}_{1},\dots,{i_{t}}_{\left|\gamma_{y}^{\min}\right|},{v_{l}}_{1},\dots,{v_{l}}_{\left|\gamma_{y}^{\min}\right|}\right\}$ of indexes of edges that are contained in $\mathcal{T}$. Then, by employing $\mathbf{F}$, program $p_{4}$ enumerates all possible minimum assembly subspaces embedded into $\mathcal{T}$ that necessarily contain every edge in $I$, thereafter choosing the $t$-th first $\Gamma^{\prime}$ that appears in this enumeration. Finally, it uses $\Gamma^{\prime}$ as input to program $p_{3}$ from the proof of Theorem 5 to return the final object $y^{\prime}$. This describes the decoding procedure from $S_{AT}\left(y\right)$. In order to construct a program $p^{\prime}_{5}$ that encodes a minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ into a $S_{AT}\left(y\right)$, one just needs to note that the codewords for $t\right>$, $\left|\gamma_{y}^{\max}\right|\right>$, $B_{\Gamma}\right>$, $v_{B}\right>$, $\gamma_{y}^{\min}\right>$, and $y\right>$ are always retrievable by an algorithm given ${\Gamma^{*}}_{y}$ as input. From these, program $p^{\prime}_{5}$ can use the same procedures that $p_{5}$ ran in order to construct $\mathcal{T}$ and enumerate its possible minimum rooted assembly subspaces. Then, it extracts $I$ from $\mathcal{T}$ by matching those that correspond to the edges and their respective labels in $\gamma_{y}^{\min}$. Next, program $p^{\prime}_{5}$ can use the same procedures that $p_{5}$ ran in order to enumerate the minimum rooted assembly subspaces embedded into $\mathcal{T}$ which contains every edge in $I$ so as to find the index $t$ that corresponds to ${\Gamma^{*}}_{y}$ in this enumeration. Finally, the program returns the sequence $S_{AT}\left(y\right)$ as output. ∎
    Proof of Theorem 9.
    With the purpose of demonstrating that $\Gamma^{\prime}={\Gamma^{*}}_{y}$, one can exploit the minimality of ${\Gamma^{*}}_{y}$ and $\gamma_{y}^{\min}$ and the maximality of $\gamma_{y}^{\max}$. First, note that there is only a finite number of rooted assembly spaces whose most distant vertices belong to $B_{\Gamma}$, which is also finite and fixed given $\Gamma$, and are separated by $\left|\gamma_{y}^{\max}\right|$ hops. From any rooted assembly space one of whose longest rooted paths is $\gamma_{y}^{\max}$, in order to construct $\mathcal{T}$, an algorithm can minimise the number of sufficient and necessary new vertices that needs to be added to $\gamma_{y}^{\max}$ by the following procedure: assign an index $j\in\mathbb{N}$ to each of the edges in $\gamma_{y}^{\max}$ by starting from the closest ones to the farthest ones from the initial vertex, which is the object in the basis set $B_{\Gamma}$, such that $1\leq j\leq\left|\gamma_{y}^{\max}\right|$; for the label of the edge $j=1$ in $\gamma_{y}^{\max}$, one only requires at most one extra vertex, and this vertex can only belong to $B_{\Gamma}$ because otherwise, since there would need to be an additional edge connecting this new vertex to the target vertex of $j=1$, it would contradict the maximality of $\gamma_{y}^{\max}$; the same analogously holds for each arbitrary $j+1\leq\left|\gamma_{y}^{\max}\right|$, so that labelling this edge $j+1$ with a vertex only requires at most one extra new vertex, and this vertex can only be at level $\leq j$, i.e., it can only belong to the set of vertices that are separated from the terminal vertex $y$ by at least $j+1$ hops; and for each new vertex created at level $j$ one would only need at most $2$ new vertices at level $\leq j-1$, and so on. Thus, by induction, one will have it that
    $\sum\limits_{i=0}^{\left|\gamma_{y}^{\max}\right|}2^{i}=2^{\left|\gamma_{y}^{\max}\right|+2}$ (12)
    is an upper bound for the maximum number of distinct vertices that a minimum rooted assembly subspace (whose $\gamma_{y}^{\max}$ is one of its longest rooted paths) can have. Following a similar inductive argument upon the maximality of $\gamma_{y}^{\max}$, one concludes that
    $2^{3\,\left|\gamma_{y}^{\max}\right|+6}$ (13)
    is an upper bound for the maximum number of distinct (multi)edges that a minimum rooted assembly subspace can have. Hence, one only needs at most
    $\mathbf{O}\left(\log\left(2^{3\,\left|\gamma_{y}^{\max}\right|+6}\right)\right)=\mathbf{O}\left(3\,\left|\gamma_{y}^{\max}\right|+6\right)$ (14)
    bits to encode either each $i_{t}$ or each $v_{l}$. (Notice that there may be tighter upper bounds than those in Equations (12) and (13), but for the purposes of Theorem 9 they suffice). Then, by hypothesis and construction, we will have it that ${\Gamma^{*}}_{y}$ can be embedded into $\mathcal{T}$, and as a consequence there is a $t\in\mathbb{N}$, where
    $1\leq t\leq\left(2^{2^{3\,\left|\gamma_{y}^{\max}\right|+6}}\right)\,\left|B_{\Gamma}\right|!$ (15)
    holds, such that $t$ corresponds to this exact structure given by ${\Gamma^{*}}_{y}$ when program $p_{5}$ starts to enumerate all possible minimum assembly subspaces embedded into $\mathcal{T}$ that necessarily contain every edge in $I$. Therefore, one needs at most
    $\mathbf{O}\left(\log\left(\left(2^{2^{3\,\left|\gamma_{y}^{\max}\right|+6}}\right)\,\left|B_{\Gamma}\right|!\right)\right)\leq\mathbf{O}\left(2^{3\,\left|\gamma_{y}^{\max}\right|+6}+\left|B_{\Gamma}\right|\log\left(\left|B_{\Gamma}\right|\right)\right)$ (16)
    bits to encode this $t$. By knowing the exact DAmG structure of ${\Gamma^{*}}_{y}$, and to which edge each basis vertex is connected, an algorithm can employ $\mathbf{F}$ to reconstruct every object assembled across this DAmG structure. Therefore, as demonstrated in Theorem 5 that from the set of vertices, the output $y^{\prime}$ is unique, one will have it that $y^{\prime}=y$. Now, by the definition of assembly space and the assembly index, we have it that $\left|\gamma_{y}^{\min}\right|\leq\left|\gamma_{y}^{\max}\right|\leq c_{\Gamma}\left(y\right)$. Therefore, from basic inequalities known from the theory of algorithmic complexity, for the best way to encode each term in the sequence in Equation (11), one obtains the proof of Equation (10). ∎
    In accordance with the authors in [1, 36, 52, 2, 4], Theorem 9 reveals that Assembly Theory’s fundamental intuition to classify the notion of complexity of assembled objects is exactly not only that of a compression scheme, but specifically an LZ scheme. The non-unidimensional structure of the minimum rooted assembly subspaces is one of the ways to define a scheme that encodes the assembling process itself back into a linear sequence of codewords.
    The encoding-decoding scheme behind Equation (11) translates the notion of the minimum assembling path in an assembly space playing a role in how much simpler the object can get, i.e., how much more compressible the object is. Shorter minimum paths lead to more compressed forms of the objects. The exponential growth in the LZ encoding from the assembly index is, in turn, consonant with the assembly number [2] of an ensemble (see Section A.3.4).
    In the same manner, Theorem 9 demonstrates that “simpler” minimum rooted assembly subspaces lead to more compressibility. For example, the reader is invited to note that due to the fact that $\gamma_{y}^{\max}=\gamma_{y}^{\min}$ would hold, a linear minimum rooted assembly subspace ${\Gamma^{*}}_{y}$, such as in [1, Figure 1(a)] or [52, Figure 8], results in a much tighter upper bound in Equation (10). This is because the exponential dependence on the assembly index to encode all possible DAG-like structures would drop: that is, $1\leq t\leq\mathbf{O}\left(\left|\gamma_{y}^{\max}\right|^{2}\,\left|B_{\Gamma}\right|\right)$ will hold instead of Equation (15), which in turn will be subsumed in the rightmost term of Equation (10) where one already has a quadratic dependence on $\left|\gamma_{y}^{\max}\right|$ or $c_{\Gamma}\left(y\right)$.
    Once a less linear assembling structure unfolds, as $\gamma_{y}^{\max}$ begins to differ more from $\gamma_{y}^{\min}$, the encoding or decoding algorithms in Theorem 9 will have to enumerate more distinct possible minimum rooted assembly spaces, and as a consequence it will increase the upper bound for $\left|S_{AT}\left(y\right)\right|$, which also renders the encoding and decoding process slower. Thus, the more complex the DAG-like structure of ${\Gamma^{*}}_{y}$ is, the more distinct DAG-like structures are possible, and therefore the closer the upper bound becomes to the exponential dependence. Hence, Equation (10) would render the object (or assembling process) less compressible, which demonstrates the claims underlying Assembly Theory that the authors in [1, 36, 52, 2, 4] originally intended to grasp with the assembly index but that in fact we demonstrated with a compression scheme. These results show that the notion of how complex an assembling process needs to be in order to construct an object is indeed equivalent to how much less compressible (by a LZ scheme) this process is.
    Further, as in Section A.3.4, one can demonstrate that any measure defined upon the assembly index (so that it increases when the assembly index(es) also increases), such as the assembly number [2], also defines an encoding and a compression method.
    A.3.4 The assembly number defines an encoding and a compression scheme
    As proposed in [2], the assembly number $A\#\left(X\right)=\sum\limits_{i=1}^{N}e^{a_{i}}\left(\frac{n_{i}-1}{N_{T}}\right)$ is intended to measure the amount of selection necessary to produce the ensemble $X$, i.e., the presence of constraints or biases in the underlying generative processes (e.g., the environment in which the objects were assembled) that set the conditions for the appearance of the objects. A higher value of $A\#\left(X\right)$ would mean that more “selective forces” were put into play as environmental constraints or biases in order to allow or generate a higher concentration of high-assembly-index elements. Otherwise, these high-assembly-index objects would not occur as often in the ensemble. This is because in stochastically random scenarios the probability rapidly decreases as the assembly index increases.
    The value of $A\#$ is calculated from the parameters $N_{T}$, $N$, $n_{i}$, $a_{i}$, which are retrievable by an algorithm from a given ensemble $X$ whose number of copies of each unique object is greater than or equal to $1$. The constant $N_{T}$ (i.e., the total number of objects in the ensemble $X$) allows one to compare the assembly number between distinct ensembles whose sizes differ.
    Notice that each term $e^{a_{i}}$ plays the role of weights for the respective ratio/proportion $\frac{n_{i}-1}{N_{T}}\leq 1$. In this manner, one can, for example, normalise again for ensembles whose overall summation of assembly indexes remains the same, hence only differing by the concentration/distribution of the quantity of objects with higher assembly index value. As stated in [1, 36, 52, 2], grasping this distinction in fact is the main goal that Assembly Theory aims at with its methods, so as to measure the presence of constraints or biases that turn otherwise unlikely objects into objects that have occurred more often. For any two ensembles $X$ and $X^{\prime}$ with the same total number $N_{T}$ of objects and whose overall summation $\sum\limits_{i=1}^{N}e^{a_{i}}$ of assembly indexes is also the same, the ensemble with the higher concentration of high-assembly-index objects (with respect to the concentration of these objects in the other ensemble) should score higher on the measure that assembly theory aims to develop. This prevents a (n equal-size) comparison between a biased sampling and another sampling in which the former was taken under distinct conditions in the latter. In this case, without normalising by the summation $\sum\limits_{i=1}^{N}e^{a_{i}}$, one could have the first one scoring a higher assembly number than the second one, but only because of the set of unique objects, and not because of the distribution of higher-assembly-index objects with respect to lower-assembly-index objects, which was what Assembly Theory intended in the first place.333 In order to avoid the presence of false positives due to abiotic mimic [62, 63], an even better control group for excluding sampling condition biases in experiments could be achieved by only comparing ensembles with not only equal size and equal total summation of assembly indexes but also the same set of unique objects. Notice that the forthcoming results in Theorem 10 with regards to assembly number hold analogously for any further normalisation than the one already given by Equation (17). For example, to tackle this issue, the value given by the measure
    $A^{\prime}\#\left(X\right)=\frac{\sum\limits_{i=1}^{N}e^{a_{i}}\left(\frac{n_{i}-1}{N_{T}}\right)}{\sum\limits_{\tiny\begin{array}[]{c}X_{j}\in\mathbf{X}\\ \sum\limits_{i=1}^{N}e^{a_{i}}=\sum\limits_{i=1}^{N_{j}}e^{a_{ij}}\end{array}}\;\left(\sum\limits_{i=1}^{N_{j}}e^{a_{ij}}\left(\frac{n_{ij}-1}{{N_{T}}_{j}}\right)\right)}$ (17)
    improves on the originally proposed measure $A\#\left(X\right)$ in [2].
    In case the set of possible objects is finite and the maximum size of the ensembles investigated is also finite, then there is a finite set $\mathbf{X}$ of possible unique ensembles. In this scenario, one can always keep effecting further renormalisations upon Equation (17) to account for any possible ensemble in $\mathbf{X}$, in this manner addressing a wider range of control groups in their experiments.
    In general, for any other measure ${A^{\prime\prime}}\#\left(\cdot\right)$ that one can possibly devise such that this measure is monotonically dependent444 That is, if $A\#\left(X_{i}\right)\leq A\#\left(X_{j}\right)$, then ${A^{\prime\prime}}\#\left(X_{i}\right)\leq{A^{\prime\prime}}\#\left(X_{j}\right)$. on $A\#\left(\cdot\right)$ (e.g., by applying any form of normalisation), one can directly apply the Krafts inequalities and the source coding theorem [58], should the ensembles be indeed sampled accordingly and the number of samples sufficiently large: one constructs an optimal code that statistically compresses each ensemble $X$ in sequences (of basis elements) whose length is $C_{A}\left(X\right)$ such that
    $-\log_{\left|B_{\mathcal{S}}\right|}\left({A^{\prime\prime}}\#\left(X\right)\right)\leq C_{A}\left(X\right)\leq-\log_{\left|B_{\mathcal{S}}\right|}\left({A^{\prime\prime}}\#\left(X\right)\right)+\mathbf{O}(1)$ (18)
    is expected to hold on average, where $\left|B_{\mathcal{S}}\right|$ is the size of the finite alphabet composed of the basis elements in $B_{\mathcal{S}}$. As proposed by Assembly Theory, notice from Equation (18) that ensembles with higher assembly numbers are expected to be more compressible, and therefore they diverge more from those ensembles that are more statistically random. This demonstrates that in the case of pure stochastic processes, Assembly Theory is either a suboptimal or an optimal compression method.
    In case the set of possible ensembles can be arbitrarily large and the ensembles are not generated by stationary and ergodic sources, finding the optimal code is not as straightforward as the one that achieved Equation (18). To tackle this generalised scenario, the algorithmic complexity theory allows one to obtain the universally optimal encoding and compression methods because Assembly Theory’s methods can always be calculated by an algorithm (i.e., its methods are recursive/computable).
    Theorem 10.
    Let $\mathcal{S}$ and $\mathbf{X}$ be enumerable by an algorithm. Let ${A^{\prime\prime}}\#\left(\cdot\right)$ be an arbitrary measure (or semimeasure) that can be effectively calculated by an algorithm from the assembly number given by the function $A\#\left(\cdot\right)$. Let $\mathbf{F}^{\prime}$ be an arbitrary formal theory that contains Assembly Theory, including all the decidable procedures of the chosen method for calculating the assembly number $A\#$ and the measure ${A^{\prime\prime}}\#\left(\cdot\right)$ for a nested subspace of $\mathcal{S}$, and the program that decides whether or not the criteria for building the assembly spaces or ensembles of objects are met. Let $X\in\mathbf{X}$ be arbitrary. Then, one can encode/compress the ensemble $X$ in $C_{A}\left(X\right)$ bits such that
    $-\log_{2}\left({A^{\prime\prime}}\#\left(X\right)\right)\leq C_{A}\left(X\right)\leq-\log_{2}\left({A^{\prime\prime}}\#\left(X\right)\right)+\mathbf{O}(1)$ (19)
    and
    $\mathbf{K}\left(X\right)\leq C_{A}\left(X\right)+\mathbf{O}(1)$ (20)
    holds, where $C_{A}\left(X\right)$ denotes the size of the codeword/program for the ensemble $X$.
    Proof.
    By hypothesis we have it that there is an algorithm that can always calculate the value of ${A^{\prime\prime}}\#\left(\cdot\right)$ for any $X$. Hence, there will be another algorithm that can always calculate
    $l_{X}\coloneqq\lceil-\log_{2}\left({A^{\prime\prime}}\#\left(X\right)\right)\rceil$ (21)
    from any ensemble $X$. Then, one directly obtains that the sequence
    $\left(\left(l_{X_{1}},X_{1}\right),\dots,\left(l_{X_{i}},X_{i}\right),\dots\right)$ (22)
    is enumerable555 In fact, not only computably enumerable, but in this case also computable by an algorithm. In addition, due to the fact that ${A^{\prime\prime}}\#\left(\cdot\right)$ is a (semi)measure by hypothesis, one has it that
    $\sum\limits_{X_{i}\in\mathbf{X}}\frac{1}{2^{l_{X_{i}}}}\leq 1\text{ .}$ (23)
    Additionally, one also has it that
    $\mathbf{K}\left(\tau_{i}\right)\leq d_{i}+\mathbf{O}(1)\text{ .}$ (24)
    From AIT [11], ones knows that if an arbitrary sequence $\left(\left(d_{1},\tau_{1}\right),\dots,\left(d_{i},\tau_{i}\right),\dots\right)$ of pairs is enumerable and $\sum\limits_{i}2^{-d_{i}}\leq 1$, then one can encode each $\tau_{i}$ in a codeword of length $d_{i}$. Therefore, from basic properties and inequalities in AIT, we will have it that there is an algorithm that can compress $X$ in a program of length $C_{A}\left(X\right)$ such that Equations (19) and (20) hold. ∎
    Now, from Theorem 10, one can employ a monotone dependence on the assembly number to demonstrate that a higher assembly number indeed implies greater compressibility of the ensemble.
    Corollary 11.
    Let the conditions for Theorem 10 be met such that ${A^{\prime\prime}}\#\left(\cdot\right)$ is also strictly monotonic on the assembly number $A\#\left(\cdot\right)$. Then, there is $k_{0}\in\mathbb{N}$ such that, if $X,X^{\prime}\in\mathbf{X}$ are ensembles with $A\#\left(X\right)+k_{0}\leq A\#\left(X^{\prime}\right)$, one has it that
    $C_{A}\left(X^{\prime}\right)<C_{A}\left(X\right)$ (25)
    holds, where $C_{A}\left(Y\right)$ denotes the size of the codeword/program for an arbitrary ensemble $Y$.
    Proof.
    Since ${A^{\prime\prime}}\#\left(\cdot\right)$ is strictly monotonic on $A\#\left(\cdot\right)$, then for every $\epsilon>0$, there is a sufficiently large $\epsilon^{\prime}>0$ such that, if $A\#\left(X\right)+\epsilon^{\prime}\leq A\#\left(X^{\prime}\right)$, one has it that ${A^{\prime\prime}}\#\left(X\right)+\epsilon\leq{A^{\prime\prime}}\#\left(X^{\prime}\right)$ holds. As a consequence, one can always pick a sufficiently large $k_{0}\in\mathbb{N}$ and arbitrary ensembles $X$ and $X^{\prime}$ where $A\#\left(X\right)+k_{0}\leq A\#\left(X^{\prime}\right)$ such that
    $-\log_{2}\left({A^{\prime\prime}}\#\left(X^{\prime}\right)\right)+\mathbf{O}(1)<-\log_{2}\left({A^{\prime\prime}}\#\left(X\right)\right)$ (26)
    holds. Therefore, Equation (25) follows from the inequality in Equation (19). ∎
    As proposed by [1, 36, 52, 2, 4], our results demonstrate that ensembles [^6] Not to be conflated with the objects themselves. with higher assembly numbers [^7] Not to be conflated with the assembly indexes. are more compressible, and therefore they diverge more from those ensembles that are outcomes of perfectly random processes. A more random generative process for an ensemble (for which the assembly number scores lower) would mean that there are fewer constraints or biases playing a role in this process. In turn, high-assembly-index objects would more rarely appear (or be constructed) because they are more complex (i.e., as we have demonstrated in the previous sections, less compressible). This is what occurs in scenarios in which there is a weaker presence of top-down (or downward) causation [41] behind the possibilities or paths that lead to the construction of the objects. Thus, as demonstrated in this section, a more compressible ensemble implies that more constraints or biases played a role in generating more unlikely (i.e., more incompressible) objects more frequently than would have been the case in an environment with otherwise fewer constraints and biases (i.e., a more random or incompressible environment), hence increasing the frequency of occurrence of high-assembly-index (i.e., less compressible) objects in this environment.
    If more random abiotic processes being present in an ensemble imply a lower assembly number, then more random abiotic processes being present in an ensemble would also imply that the ensemble is more incompressible.
    Therefore, along with Section A.2, the results in this Section A.3.4 demonstrate that in the general case (even when the underlying generative process of the ensembles is unknown) assembly number $A\#\left(\cdot\right)$ is an approximation to algorithmic complexity (or algorithmic probability), but a suboptimal measure with respect to the more general methods from algorithmic information theory.
    A.4 Lack of empirical evidence in the favour of Assembly Theory and empirical evidence against it
    The experiments and analysis of organic versus non-organic molecules from mass spectral data using several complexity indexes used as positive evidence for Assembly Theory in [36] empirically demonstrates what we have proven. That the assembly index is equivalent to statistical compression, given that it was not compared to any other data type or any other existing index of the same nature. The extent of the simplicity of the experiments required to prove the utility of the assembly index and the main hypothesis of Assembly Theory makes us believe that this was a serious omission.
    Together with the theoretical results, Figure 3 from [16] shows that the evidence that AT presents as being in favour of their theory and algorithms, specifically that their assembly compression index may correspond to the way objects assemble, is flawed and invalid. The empirical evidence actually shows that other, equally plausible and valid compression schemes that are sequential, count copies, and build a repeating symbol dictionary tree, are as equally or more likely based on their discriminatory power that Assembly Theory uses as evidence in their favour.
    Figure 3: Taken from [16]. MA stands for Molecular Assembly based on the assembly index of the molecular compounds. These results were obtained even without making any optimisations to the various algorithms after a trivial flattenization procedure of the data (which preserves algorithmic complexity but can help the non-MA indexes converge faster)
    The strongest positive correlation in Fig. 3, was identified between MA and 1D-RLE coding (R= 0.9), which is one of the most basic coding schemes known in computer science since the 60s and among the most similar to the intended definition of MA, as being capable of ‘counting copies’ together with LZ that is exactly equivalent to MA or the assembly index. All indexes were applied to spectral data of the 132 molecules (114 plus 18 living extracts) from the original Assembly Theory paper [36] made available by the authors of AT. Other coding algorithms, including the Huffman coding (R = 0.896), also show a strong positive correlation with MA. The analysis further confirms our previous findings of the similarity in performance between MA and other basic compression measures (that essentially only ‘count copies’ in data, all of them converging to Entropy) in classifying living vs. non-living mass spectra signatures.